# Cosmic baryon census with fast radio bursts and gravitational waves

### v2

@@ Abstract
The cosmic baryon density fraction ($\Omega_{\rm b}$) is intrinsically correlated with the Hubble constant ($H_{0}$) through the critical density of the Universe. In the context of the decade-long $H_{0}$ tension, the significant discrepancy between early- and late-Universe measurements of $H_{0}$ implies that fixing its value or imposing an external prior could bias the baryon census. To address this concern, we construct a late-Universe probe framework that unifies fast radio bursts (FRBs) and gravitational-wave (GW) standard sirens, which can respectively resolve the “missing baryon” problem and the $H_{0}$ tension through their dispersion measures and absolute luminosity distances. By combining $104$ localized FRBs with $47$ GW events, we obtain an $H_{0}$-free measurement of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), in concordance with early-Universe observations of CMB + BBN. Although the current precision ($\sim 13\%$) is limited by sample size, the growing detections of both FRBs and GWs will make their synergy a powerful probe of low-redshift cosmology.
I Introduction
The concordance cosmology predicts consistency among independent probes across all epochs Ostriker:1995su. However, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe Cooke:2017cwo, about $30\%$ baryons remain undetected in the local Universe Shull:2011aa. These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” Fukugita:1997bi; cen1999baryons.
@@ I Introduction
Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin lorimer2007; thornton2013pop, which can be broadly classified into repeaters and non-repeaters 2016Natur.531..202S; Sun2025a; Sun2025b. Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) mcquinn2013locating. Thus, FRBs with known redshifts derived from precise host galaxy localization can be used to trace the missing baryons Deng:2013aga; connor2025gas. Ref. Macquart:2020lln (hereafter Macquart20) first employed $5$ localized FRBs to estimate the present-day cosmic baryon density of $\Omega_{\rm b}=0.051^{+0.021}_{-0.020}$ at the $95\%$ confidence level, which agrees with the joint CMB and BBN result and likely resolves the local baryon problem. Ref. Yang:2022ftm (hereafter Yang22) used an updated sample of $22$ localized FRBs and obtained $\Omega_{\rm b}=0.0490_{-0.0033}^{+0.0036}$ at the $1\sigma$ level, reinforcing the evidence for a complete cosmic baryon census. With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe Zhou:2014yta; Gao:2014iva; dai2023ska; Zhang:2023gye; see Refs. Bhandari:2021thi, Wu:2024iyu, and Glowacki:2024cgu for recent reviews.
@@
However, FRB cosmology based on DM is hindered by parameter degeneracies, which result in two main sources of systematic uncertainties Kumar:2019qhc. The first is from extracting the IGM contribution ($\mathrm{DM}_{\rm IGM}$) from the total DM budget (see Eq. (1)). This is mainly due to the difficulty in modeling both the host contribution ($\mathrm{DM}_{\rm host}$) and the LoS inhomogeneity in the IGM. Macquart20 modeled their probability distributions and Refs. Zhang:2020mgq; Zhang:2020xoc further employed the IllustrisTNG simulations to derive priors on these empirical DM model parameters for different FRB populations, which effectively reduce biases induced by host galaxy diversity and IGM fluctuations. However, the second uncertainty from the analysis of the extracted $\mathrm{DM}_{\rm IGM}$ remains challenging to control. This is because its redshift relation (known as the Macquart relation) exhibits intrinsic degeneracies among $\Omega_{\rm b}$, the Hubble constant ($H_{0}$), and the baryon fraction in diffuse ionized gas ($f_{\rm d}$) (see Eq. (3)). Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature Li:2020qei; Qiang:2020vta; Wang:2022ami; Liu:2025fdf. $\Omega_{\rm b}$ is intrinsically linked to $H_{0}$ through the critical density of the Universe. Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB 2020Planck (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) Riess:2021jrx) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the $H_{0}$ tension” Guo:2018ans; DiValentino:2021izs; Abdalla:2022yfr; Hu:2023jqc; CosmoVerseNetwork:2025alb. This inconsistency would introduce systematic uncertainties into the baryon census using FRB samples. It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 67677474 km s-1  Mpc-1, which provides a reasonable strategy for handling the second uncertainty given the $H_{0}$ tension (note that many studies also constrain $H_{0}$ by adopting early-Universe priors on $\Omega_{\rm b}h^{2}$; see, e.g., Ref. Hagstotz:2021jzu). Nevertheless, any external assumption on $H_{0}$ could bias the baryon estimate, motivating the need for a direct inference of $\Omega_{\rm b}$ that is entirely untethered to the Hubble tension, e.g., jointly analyzing FRBs and datasets that independently measure $H_{0}$ in the late Universe.
@@
As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ Schutz:1986gp. This primarily arises from the fact that the amplitudes of GW waveforms are inversely proportional to the luminosity distances ($D_{\rm L}$) to the sources, allowing for a direct measurement of $D_{\rm L}$ through waveform analysis. Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known Dalal:2006qt; Cutler:2009qv; Zhao:2010sz; Cai:2016sby; LIGOScientific:2017adf; LIGOScientific:2017vwq; LIGOScientific:2017zic; Chen:2017rfc; Cai:2017aea; Wang:2018lun; Zhang:2019loq; Zhang:2018byx; Zhang:2019ylr; Wang:2019tto; Zhao:2019gyk; Gray:2019ksv; Jin:2020hmc; Yu:2020vyy; Mastrogiovanni:2021wsd; KAGRA:2021duu; Wang:2021srv; Zhu:2021aat; Zhu:2021bpp; Wu:2022dgy; Muttoni:2023prw; Yun:2023ygz; Yu:2023ico; Han:2023exn; Feng:2024lzh; Song:2022siz; Jin:2023sfc; Jin:2023tou; Li:2023gtu; Dong:2024bvw; Xiao:2024nmi; Zhu:2024qpp; Zheng:2024mbo; Han:2025fii; Du:2025odq; Song:2025bio; Jin:2025dvf; Dong:2025ikq. During the first three observing runs of the LIGO, Virgo, and KAGRA (LVK) collaboration, approximately 100 GW events were detected and compiled into the GWTC-3 catalog. Using 47 GW standard sirens from GWTC-3, Ref. LIGOScientific:2021aug achieved a $\sim 10\%$ precision on $H_{0}$, and Ref. Song:2025ddm further combined these sirens with an additional strong-lensing system to obtain a model-independent $\sim 8\%$ constraint. More recently, with the release of the LVK O4a data set, over 300 GW events have been reported, and by analyzing 142 of them while allowing the population-model parameters of both binary black holes (BBHs) and binary neutron stars (BNSs) to vary freely, the LVK collaboration has obtained a $\sim 14\%$ constraint on $H_{0}$ LIGOScientific:2025jau. Therefore, the $H_{0}$ information uniquely provided by GW data enables a direct estimate of $\Omega_{\rm b}$ disentangled from $H_{0}$ in the Macquart relation, which offers new insight into resolving systematic effects in FRB cosmology. Unlike previous studies that combined FRBs with traditional observations such as SNe Jaroszynski:2018vgh; gjz2023 or CMB Zhao:2020ole; Qiu:2021cww, we highlight that combining with GWs inherently involves calibration-free distances and avoids reliance on early-time physics, thereby eliminating the bias induced by the $H_{0}$ tension.
@@
In this work, we present the first $H_{0}$-free baryon census at low redshift, within a late-Universe probe framework that unifies $104$ localized FRBs and $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$. To control model-dependent systematics, we impose no FRB population model and fully marginalize over all DM parameters [^1] In this context, we refer to the DM parameters as ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$), while the empirical DM model parameters refer specifically to ($F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$). The parameters are defined in Sec. II.1., while on the GW side we adopt a population model for the main analysis and additionally release its parameters to assess robustness. Note that we do not assume any FRB/GW associations (see Refs. Wei:2018cgd and Qiang:2024erm for studies based on such associations).
This work is organized as follows. In Sec. II, we describe the datasets and likelihood construction for both FRBs and GWs. Results and related discussions are presented in Sec. III, and we conclude in Sec. IV.
II Data and methodology
II.1 Localized FRB data
In this section, we first present the likelihood function for localized FRBs based on DM analysis following Macquart20, then the FRB sample used in this work.
@@ II.1 Localized FRB data
where $\Delta\equiv{\rm{DM}}_{\rm{IGM}}/\left\langle{\rm{DM}}_{{{\rm{IGM}}}}\right\rangle$. $\alpha$ and $\beta$, which describe the inner gas density profile of halos, are best fitted as $\alpha=3$ and $\beta=3$ according to hydrodynamic simulations. The effective standard deviation $\sigma_{\mathrm{IGM}}=Fz^{-1/2}$ is parameterized by Macquart20, where $F$ is the parameter reflecting the galactic feedback. Here, $A$ is a normalization constant, and $C_{0}$ is chosen to ensure that the mean of this distribution is unity. Note that we include the factor $1/\langle{\rm DM_{\rm IGM}}\rangle$ suggested by Ref.  Zhang:2025wif when normalizing Eq. (4).
@@
For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs. (2), (4), and (II.1):
@@
We also consider a scenario in which the empirical DM model parameters are fixed, though our main results treat them as free. Specifically, we adopt values calibrated from cosmological hydrodynamical simulations, notably the IllustrisTNG simulation, which provides redshift-dependent distributions for both $P_{\rm IGM}(\mathrm{DM}_{\rm IGM})$ Zhang:2020xoc and $P_{\rm host}(\mathrm{DM}_{\rm host})$ Zhang:2020mgq. The parameters $A(z)$, $C_{0}(z)$, and $\sigma_{\rm IGM}(z)$ in Eq. (4), as well as $\mu_{\rm host}(z)$ and $\sigma_{\rm host}(z)$ in Eq. (II.1), are interpolated from best-fit simulation results across redshift. Note that $\mu_{\rm host}$ and $\sigma_{\rm host}$ are further fixed based on the host galaxy types (see the seventh column of Table LABEL:tab:localized_FRBs).
@@
Finally, we select a sample of $104$ localized FRBs from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ connor2025gas; (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments. The final sample is listed in Appendix B (see Tables LABEL:tab:localized_FRBs).
II.2 GW standard siren data
@@
We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3222https://gwosc.org/GWTC-3, covering a redshift range of approximately $z<0.8$. The sample consists of 42 BBHs, three neutron star-black holes (NSBHs), and two BNSs. We do not consider GW events detected in the LVK O4a run, mainly because during this run, Virgo is not online, and all GW events are detected by two LIGOs, resulting in relatively poorer sky localizations and luminosity distance measurements LIGOScientific:2025jau.
@@
We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart LIGOScientific:2017adf; LIGOScientific:2017vwq; LIGOScientific:2017zic. For GW events without EM counterparts, we infer their redshifts through the dark siren method Schutz:1986gp; Mastrogiovanni:2023emh; Gray:2023wgj and incorporate the GLADE+ catalog Dalya:2021ewn. In the dark siren method, assuming the GW sources all originate from galaxies, we can cross-match sky localizations of GW events with galaxies collected in galaxy catalogs, and identify a serious potential host galaxy. These potential host galaxies provide redshift priors for GW events, which help to infer cosmological parameters in a hierarchical Bayesian framework. Furthermore, GW population models can help break degeneracies of masses and redshifts in GW detections, thereby providing additional redshift information Taylor:2011fs; Ezquiaga:2022zkx; Mastrogiovanni:2023emh; Gray:2023wgj. In this paper, we use the public pipeline ICAROGW Mastrogiovanni:2023zbw to compute the GW likelihood function, denoted as $\mathcal{L}_{\rm GW}$. The explicit formulas can be found in Refs Mastrogiovanni:2023emh; Gray:2023wgj.
@@
We adopt the same population model assumptions as Ref. LIGOScientific:2021aug, but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the GWTC-3 data. Considering additional models would increase computational cost and lies beyond the scope of this work. For BNSs, we assume that neutron star masses are uniformly distributed in the range of 1–3 $M_{\odot}$. In the case of NSBH systems, the black hole component is drawn from the same distribution as the primary black holes in BBHs, while the neutron star component follows the mass distribution assumed for BNSs. The redshift distribution of GW sources is modeled using a phenomenological prescription that traces the cosmic star formation rate. Consistent with Ref. LIGOScientific:2021aug, we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution. These effects are currently subdominant compared to the statistical uncertainties present in the existing GW observations.
II.3 Joint analysis of FRB and GW data
Figure 1: Normalized posterior distributions of $\Omega_{\rm b}$~from 104 FRBs (orange) and 104 FRBs + 47 GWs (indigo). The median and $1\sigma$ credible interval from FRB + GW is $\Omega_{\rm b}=0.0488\pm 0.0064$, marginalized over $f_{\rm d}$ and empirical DM model parameters. The errorbars in different colors and linestyles are shown under different assumptions: FRB + $H_{0}$ (Planck18 prior) and FRB + $H_{0}$ (SH0ES prior) adopt Gaussian priors on $H_{0}$ from Planck18 2020Planck and SH0ES Riess:2021jrx, respectively; FRB (TNG) + GW and FRB (TNG + $f_{\rm d}(z)$) + GW adopt priors on empirical DM model parameters from the IllustrisTNG simulation Zhang:2020mgq; Zhang:2020xoc and on $f_{\rm d}$ from the redshift-dependent form of Macquart20 Macquart:2020lln, respectively. The blue, gray, and red shaded bands show the results from CMB + BBN Cooke:2017cwo, Macquart20, and Yang22 Yang:2022ftm, respectively. Note that Yang22 adopts the TNG prior, a flat prior on $H_{0}$ $\in$ $\mathcal{U}(67,73)$ km s-1 Mpc-1, and $f_{\rm d}=0.82$. All results are at the $1\sigma$ level.
@@ II.3 Joint analysis of FRB and GW data
We jointly constrain cosmological and FRB-related DM parameters, and adopt flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ Yang:2022ftm, $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^3] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs. Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ connor2025gas. $F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ Macquart:2020lln. For the GW population model parameters, we consider two scenarios. In the first, following Ref. LIGOScientific:2021aug, we fix all population-model parameters to their fiducial values. In the second, we allow the BBH mass-distribution parameters and the redshift-distribution parameters of GW sources to vary freely, adopting the same prior ranges as the spectral siren analysis in Ref. LIGOScientific:2021aug, to assess the impact of population-model uncertainties on our results. This treatment is similar to Ref. Gray:2023wgj; however, unlike Ref. LIGOScientific:2025jau, we do not free the mass-distribution parameters of neutron-star sources.
@@
Table 1: $1\sigma$ constraints on cosmological parameters $\Omega_{\rm b}$ and $H_{0}$. The results are derived from $104$ localized FRBs alone or in combination with $47$ GW events, under different assumptions. For all cases, the $1\sigma$ constraints on DM parameters are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$. Unless otherwise specified, the FRB + GW case assumes a flat prior of $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ and fixed GW population parameters. Note that $H_{0}$ is in units of km s-1  Mpc-1.

### v3

@@ Abstract
The cosmic baryon density fraction ($\Omega_{\rm b}$) is intrinsically correlated with the Hubble constant ($H_{0}$) through the critical density of the Universe. In the context of the decade-long $H_{0}$ tension, the significant discrepancy between early- and late-Universe measurements of $H_{0}$ implies that fixing its value or imposing an external prior could bias the baryon census. To address this concern, we construct a late-Universe probe framework that unifies fast radio bursts (FRBs) and gravitational-wave (GW) standard sirens, which can respectively resolve the “missing baryon” problem and the $H_{0}$ tension through their dispersion measures (DMs) and absolute luminosity distances. By combining $104$ localized FRBs with $47$ GW events, we obtain an $H_{0}$-free measurement of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), in concordance with early-Universe observations of CMB + BBN. The result is tightly anchored by GW-inferred $H_{0}$ through the strong $\Omega_{\rm b}$-$H_{0}$ degeneracy. Although the current precision ($\sim 13\%$) is limited by sample size, the growing detections of both FRBs and GWs will make their synergy a powerful probe of low-redshift cosmology.
Introduction.
@@
The concordance cosmology predicts consistency among independent probes across all epochs Ostriker and Steinhardt (1995). Yet, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe Cooke et al. (2018), about $30\%$ baryons remain undetected in the local Universe Shull et al. (2012). These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” Fukugita et al. (1998); Cen and Ostriker (1999). Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin Lorimer et al. (2007); Thornton et al. (2013), which can be broadly classified into repeaters and non-repeaters Spitler et al. (2016); Sun et al. (2025a, b). Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) McQuinn (2013). Localized FRBs with measured redshifts thus provide a direct probe of the missing baryons Deng and Zhang (2014); Connor et al. (2025). Using 5 and 22 localized FRBs, Refs. Macquart et al. (2020); Yang et al. (2022) (hereafter Macquart20 and Yang22) inferred $\Omega_{\rm b}$ values consistent with CMB + BBN, suggesting a nearly complete baryon census. With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe to, e.g., probe dark energy Zhou et al. (2014); Gao et al. (2014); Walters et al. (2018); Dai (2023); Zhang et al. (2023), the Hubble constant ($H_{0}$) Li et al. (2018); Hagstotz et al. (2022); Wu et al. (2022); James et al. (2022); Liu et al. (2023); Zhao et al. (2022); Kalita et al. (2025); Fortunato et al. (2025); Zhang et al. (2025a); Wang et al. (2025a); Gao et al. (2025a); Zhang and Zhang (2025); Gao et al. (2025b); Zhuge et al. (2026); Xu et al. (2025); Jia et al. (2025); Wu et al. (2026), as well as baryonic feedback Sharma et al. (2026); Reischke and Hagstotz (2025), and the large-scale structure Hussaini et al. (2025); Wang et al. (2025b); see Refs. Bhandari and Flynn (2021); Petroff et al. (2022); Xiao et al. (2021); Zhang (2023); Wu and Wang (2024); Glowacki and Lee (2025) for recent reviews.
@@
However, FRB cosmology based on DM is seriously hindered by parameter degeneracies, which result in two main sources of systematic uncertainties Kumar and Linder (2019). The first is from extracting the IGM contribution ($\mathrm{DM}_{\rm IGM}$) from the total DM budget (see Eq. (1)). This is mainly due to the difficulty in modeling both the host contribution ($\mathrm{DM}_{\rm host}$) and the LoS inhomogeneity in the IGM. Macquart20 modeled their probability distributions and Refs. Zhang et al. (2020, 2021) further employed the IllustrisTNG (TNG) simulations to derive priors on these empirical model parameters for different FRB populations, which could help mitigate biases from host diversity and IGM fluctuations. However, the second, and more fundamental, uncertainty from the analysis of the extracted $\mathrm{DM}_{\rm IGM}$ remains challenging to control. This is because its redshift relation (known as the Macquart relation) exhibits intrinsic degeneracies among $\Omega_{\rm b}$, $H_{0}$, and the baryon fraction in diffuse ionized gas ($f_{\rm d}$) (see Eq. (3)). Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature Jaroszynski (2019); Li et al. (2020); Qiang and Wei (2020); Zhao et al. (2020a); Qiu et al. (2022); Wang and Wei (2023); Gao et al. (2023); Liu et al. (2025). On the other hand, $H_{0}$ is coupled to $\Omega_{\rm b}$ through the critical density of the Universe. Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB Aghanim et al. (2020) (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) Riess et al. (2022)) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the $H_{0}$ tension” Guo et al. (2019); Di Valentino et al. (2021); Abdalla et al. (2022); Hu and Wang (2023); Di Valentino et al. (2025). This inconsistency would introduce systematic uncertainties into the baryon census using FRB samples. It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 67677474 km s-1  Mpc-1, which could unavoidably bias the baryon estimate in the context of the Hubble tension and thereby motivate a direct inference of $\Omega_{\rm b}$, e.g., jointly analyzing FRBs and datasets that independently measure $H_{0}$ in the late Universe.
@@
As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ Schutz (1986). This primarily arises from the fact that the amplitudes of GW waveforms are inversely proportional to the luminosity distances ($D_{\rm L}$) to the sources, allowing for a direct measurement of $D_{\rm L}$ through waveform analysis. Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known Dalal et al. (2006); Cutler and Holz (2009); Zhao et al. (2011); Cai and Yang (2017); Abbott et al. (2017a, b, c); Chen et al. (2018); Cai et al. (2018); Wang et al. (2018); Zhang et al. (2019a, b); Zhang (2019); Wang et al. (2020); Zhao et al. (2020b); Gray et al. (2020); Jin et al. (2020); Yu et al. (2020); Mastrogiovanni et al. (2021); Abbott et al. (2023a); Wang et al. (2022); Zhu et al. (2022a, b); Wu et al. (2023); Muttoni et al. (2023); Yun et al. (2023); Yu et al. (2024); Han et al. (2024); Feng et al. (2024); Song et al. (2024); Jin et al. (2024a, b); Li et al. (2024); Dong et al. (2025a); Xiao et al. (2025); Zhu et al. (2024); Zheng et al. (2024); Han et al. (2026); Du et al. (2025); Song et al. (2026); Jin et al. (2026); Dong et al. (2025b). During the first three observing runs of the LIGO, Virgo, and KAGRA (LVK) collaboration, approximately 100 GW events were detected and compiled into the GWTC-3 catalog. Using 47 GW standard sirens from GWTC-3, Ref. Abbott et al. (2023b) achieved a $\sim 10\%$ precision on $H_{0}$, and Ref. Song et al. (2025) further combined these sirens with an additional strong-lensing system to obtain a model-independent $\sim 8\%$ constraint. More recently, with the release of the LVK O4a data set, over 300 GW events have been reported, and by analyzing 142 of them while varying the population-model parameters of both binary black holes (BBHs) and binary neutron stars (BNSs), the LVK collaboration has obtained a $\sim 14\%$ constraint on $H_{0}$ Abac et al. (2025). Therefore, the $H_{0}$ information uniquely provided by GW data enables a direct estimate of $\Omega_{\rm b}$ disentangled from $H_{0}$ in the Macquart relation, which offers new insight into resolving systematic effects in FRB cosmology.
@@
In this Letter, we present the first $H_{0}$-free baryon census at low redshift, within a late-Universe probe framework that unifies $104$ localized FRBs and $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$. Note that we do not assume any FRB/GW associations (see Refs. Wei et al. (2018); Qiang et al. (2025); Jahns-Schindler and Spitler (2025) for studies based on such associations).
Localized FRBs.
@@
where $\Delta\equiv{\rm{DM}}_{\rm{IGM}}/\left\langle{\rm{DM}}_{{{\rm{IGM}}}}\right\rangle$. $\alpha$ and $\beta$, which describe the inner gas density profile of halos, are best fitted as $\alpha=3$ and $\beta=3$ according to hydrodynamic simulations. The effective standard deviation $\sigma_{\mathrm{IGM}}=Fz^{-1/2}$ is parameterized by Macquart20, where $F$ is the parameter reflecting the galactic feedback. Here, $A$ is a normalization constant, and $C_{0}$ is chosen to ensure that the mean of this distribution is unity. Note that we include the factor $1/\langle{\rm DM_{\rm IGM}}\rangle$ suggested by Ref.  Zhang et al. (2025b) when normalizing Eq. (4).
@@
For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs. (2), (4), and (Localized FRBs.):
@@
Finally, we select a sample of $104$ localized FRBs (see the Supplemental Material SM (see also Refs. Chatterjee et al. (2017); Tendulkar et al. (2017); Bhandari et al. (2022); Bannister et al. (2019); Prochaska et al. (2019); Bhardwaj et al. (2024); Bhandari et al. (2020); Ibik et al. (2024); Michilli et al. (2023); Ravi et al. (2019); Chittidi et al. (2021); Heintz et al. (2020); Law et al. (2020); Shin et al. (2025); Ravi et al. (2022); Shannon et al. (2025); Caleb et al. (2023); Cassanelli et al. (2024); Gordon et al. (2023); Sharma et al. (2024); Law et al. (2024); Li et al. (2026); Ryder et al. (2023); Rajwade et al. (2024); Anna-Thomas et al. (2025); Hanmer et al. (2025); Connor et al. (2025); Eftekhari et al. (2025); Amiri et al. (2025); Tian et al. (2024) therein). The dataset is selected from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ Connor et al. (2025); (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments.
GW standard sirens.
@@
We use 47 GW standard sirens with SNRs greater than 11 and an inverse false alarm rate exceeding 4 yr from GWTC-3111https://gwosc.org/GWTC-3, covering a redshift range of approximately $z<0.8$. The sample consists of 42 BBHs, three neutron star-black holes (NSBHs), and two BNSs. We do not consider GW events detected in the LVK O4a run, mainly because during this run, Virgo is not online, and all GW events are detected by two LIGOs, resulting in relatively poorer sky localizations and luminosity distance measurements Abac et al. (2025).
@@
We obtain the redshift of GW170817 from the electromagnetic (EM) counterpart Abbott et al. (2017a, b, c). For GW events without EM counterparts, we infer their redshifts through the dark siren method Schutz (1986); Mastrogiovanni et al. (2023); Gray et al. (2023) and incorporate the GLADE+ catalog Dálya et al. (2022). In the dark siren method, assuming the GW sources all originate from galaxies, we can cross-match sky localizations of GW events with galaxies collected in galaxy catalogs, and identify a serious potential host galaxy. These potential host galaxies provide redshift priors for GW events, which help to infer cosmological parameters in a hierarchical Bayesian framework. Furthermore, GW population models can help break degeneracies of masses and redshifts in GW detections, thereby providing additional redshift information Taylor et al. (2012); Ezquiaga and Holz (2022); Mastrogiovanni et al. (2023); Gray et al. (2023). In this paper, we use the public pipeline ICAROGW Mastrogiovanni et al. (2024) to compute the GW likelihood function, denoted as $\mathcal{L}_{\rm GW}$. The explicit formulas can be found in Refs Mastrogiovanni et al. (2023); Gray et al. (2023).
@@
We adopt the same population model assumptions as Ref. Abbott et al. (2023b), but restrict our analysis to the Power Law + Peak BBH mass model, which is best supported by the GWTC-3 data. Considering additional models would increase computational cost and lies beyond the scope of this work. For BNSs, we assume that neutron star masses are uniformly distributed in the range of 131-3 $M_{\odot}$. In the case of NSBH systems, the black hole component is drawn from the same distribution as the primary black holes in BBHs, while the neutron star component follows the mass distribution assumed for BNSs. The redshift distribution of GW sources is modeled using a phenomenological prescription that traces the cosmic star formation rate. Consistent with Ref. Abbott et al. (2023b), we disregard potential redshift evolution in the BBH mass distribution as well as variations in the GW detection rate arising from the spin distribution. These effects are currently subdominant compared to the statistical uncertainties present in the existing GW observations.
Joint analysis of FRB and GW data.
@@
We jointly constrain cosmological and all DM parameters with flat priors for the following seven parameters: $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ Yang et al. (2022), $\Omega_{\rm b}\in\mathcal{U}(0.01,0.10)$, $H_{0}\in\mathcal{U}(20,140)~\rm km\ s^{-1}\ Mpc^{-1}$, $f_{\rm d}\in\mathcal{U}(0.84,0.96)$, [^2] This prior on $f_{\rm d}$ is informed by observations of the baryon fraction in stars ($f_{*}$) and cold gas ($f_{\rm cold}$), independent of FRBs. Assuming a conservative flat prior for $f_{*}$ extended by a factor of five, the diffuse baryon fraction is inferred via $f_{\rm d}=1-f_{*}-f_{\rm cold}$ Connor et al. (2025). $F\in\mathcal{U}(0.011,1.0)$, $\mu_{\rm host}\in\mathcal{U}(3.0,5.3)$, and $\sigma_{\rm host}\in\mathcal{U}(0.2,2.0)$ Macquart et al. (2020). To test the robustness of our analysis, we examine the effect of prior assumptions on our results. For the $\Omega_{\rm m}$ prior, we relax it to a broad range $\mathcal{U}(0,1)$ to assess its impact on the constraints. For the GW population-model parameters, following Ref. Abbott et al. (2023b), we first fix all parameters to their fiducial values and then, vary the BBH mass- and redshift-distribution parameters of GW sources, adopting the same prior ranges as in the spectral-siren analysis of Ref. Abbott et al. (2023b), to assess the impact of population-model uncertainties on our results. This treatment is similar to Ref. Gray et al. (2023); however, unlike Ref. Abac et al. (2025), we do not free the mass-distribution parameters of neutron-star sources.
@@
Table 1: $1\sigma$ constraints on cosmological parameters $\Omega_{\rm b}$ and $H_{0}$. The results are derived from 104104 localized FRBs alone or in combination with $47$ GW events, under different assumptions. Unless otherwise specified, the FRB + GW case assumes a flat prior of $\Omega_{\rm m}\in\mathcal{U}(0.28,0.35)$ and fixed GW population parameters. Note that $H_{0}$ is in units of km s-1  Mpc-1.